What is the amount of hydrogen fluoride you get from the action of 6.4g caf2 with excess sulfuric acid concentrated (h2so4)
1
Expert's answer
2020-05-21T13:53:47-0400

Solution.

"CaF2 + H2SO4 = CaSO4 + 2HF"

"n(CaF2) = \frac{6.4}{78.07} = 0.082 \ mol"

"n(HF) = 2 \times n(CaF2) = 0.164 \ mol"

"m(HF) = 0.164 \times 20.00 = 3.28 \ g"

Answer:

m(HF) = 3.28 g